### v1

# Flow Matching for Tabular Data Synthesis

Abstract
Synthetic data generation is an important tool for privacy-preserving data sharing. While diffusion models have set recent benchmarks, flow matching (FM) offers a promising alternative. This paper presents different ways to implement flow matching for tabular data synthesis. We provide a comprehensive empirical study that compares flow matching (FM and variational FM) with a state-of-the-art diffusion method (TabDDPM and TabSyn) in tabular data synthesis. We evaluate both the standard Optimal Transport (OT) and the Variance Preserving (VP) probability paths, and also compare deterministic and stochastic samplers – something possible when learning to generate using variational flow matching – characterising the empirical relationship between data utility and privacy risk. Our key findings reveal that flow matching, particularly TabbyFlow, outperforms diffusion baselines. Flow matching methods also achieves better performance with remarkably low function evaluations ($\leq$ 100 steps), offering a substantial computational advantage. The choice of probability path is also crucial, as using the OT path demonstrates superior performance, while VP has potential for producing synthetic data with lower disclosure risk. Lastly, our results show that making flows stochastic not only preserves marginal distributions but, in some instances, enables the generation of high utility synthetic data with reduced disclosure risk.
1 Introduction
Government bodies, particularly national statistical offices (NSOs), are confronted with a pressing challenge: how to disseminate useful tabular data while preserving individual privacy. Tabular data, which are structured data with rows and columns representing real-world entities and their attributes, are essential for decision-making in economics, healthcare, social sciences, and finance (Little2024synthetic). However, because these datasets often contain sensitive attributes, such as financial conditions or medical diagnoses, privacy protection laws such as the EU’s General Data Protection Regulation (GDPR) and Indonesia’s Personal Data Protection (PDP) Act limit access to them (ruggles2021role). This tension between data utility and confidentiality has spurred interest in methods that can generate synthetic tabular data that preserve statistical properties while reducing the risk of disclosure (nowok2016synthpop1).
Recent advances in deep generative models have provided a promising approach to tabular data synthesis. Deep learning techniques, such as Generative Adversarial Networks (goodfellow2014generative), Variational Autoencoders (kingma2013auto1), and more recently, diffusion models (dickstein2015deep; ho2020denoising), have demonstrated impressive capabilities in generating synthetic data of various forms, including images and text, as well as tabular data (xu2019modeling; kotelnikov2023tabddpm).
Diffusion models, in particular, have emerged as a powerful option for generative tasks, including the synthesis of tabular data. TabDDPM (kotelnikov2023tabddpm), a diffusion model tailored for tabular data, and other models such as TabSyn (zhang2024mixedtype) have gained attention for their ability to model intricate dependencies between variables in tabular datasets by iteratively transforming noise into structured data. The main strength of these models lies in their ability to handle the high-dimensional structure of tabular data, where the relationships between variables are often nonlinear and complex.
Another recent development is flow matching (FM) (lipman2023flowmatchinggenerativemodeling; albergo2023stochasticinterpolantsunifyingframework; liu2022rectifiedflowmarginalpreserving), which approximates the data distribution through a flow, i.e. through learning a velocity field that induces an ordinary differential equation transporting noise to data. Variational Flow Matching (VFM) (eijkelboom2024variationalflowmatchinggraph) offers a reinterpretation of flow matching as a form of variational inference over trajectories, providing an elegant solution to parameterise the flow towards any kind of distribution (e.g. also discrete or constrained). This method holds significant promise for tabular data synthesis, as it ensures that the generated output not only mirrors the statistical attributes of the original data but also captures patterns that traditional generative techniques might overlook. Beyond tabular data generation (guzman-cordero2025exponential), VFM has also seen recent success in fields like molecular generation (zaghen2025riemannianvariationalflowmatching; eijkelboom2025controlled; sakalyan2025modeling), image generation (maticsan2025purrception), and climate modeling (finn2025generative). Further discussion on related work can be found in Appendix A.
Although recent advances in tabular diffusion models (kotelnikov2023tabddpm; zhang2024mixedtype) and FM (guzman-cordero2025exponential) have shown promise, existing studies remain limited in four ways. First, they have focused primarily on open source benchmark datasets, neglecting the unique challenges of census data as a real-world representation. Second, utility and privacy risk evaluation has been restricted to computer science metrics such as ML accuracy and distance to closest record (DCR), which (i) may not align with statistical needs,(ii) are computationally intensive, and (iii) can be hard to interpret for practitioners. Third, recent flow matching studies (eijkelboom2024variationalflowmatchinggraph; guzman-cordero2025exponential; eijkelboom2025controlled) have predominantly focused on optimal transport trajectories. The potential of variance-preserving trajectories - although theoretically established in the original FM formulation (lipman2023flowmatchinggenerativemodeling) and proven effective in diffusion models (ho2020denoising) — remains underexplored. Finally, the interaction between latent vs. direct representations and deterministic vs. stochastic dynamics in FM-based tabular synthesis remains unclear, leaving open questions about which configuration produces the best synthetic data.
Research Objective. This paper’s primary goal is quantify whether flow matching can offer an efficient, high-utility, and low-risk alternative to diffusion models for mixed-type tabular data. We address the goal through four specific research questions.
  • Q1. Can FM-based models improve upon diffusion model baselines in tabular data synthesis? (Section 5.1)
  • Q2. How efficient are FM models in terms of the utility and risk of the generated output? (Section 5.2)
  • Q3. How do utility and risk characteristics of the synthetic data evolve with integration time? (Section 5.3)
  • Q4. Can stochastic sampling improve synthesis quality relative to deterministic sampling? (Section 5.4)
Contributions. We explore flow matching techniques for tabular data synthesis through four implementations: (1) learning the distribution of latent variables using regular flow matching, inspired by TabSyn (zhang2024mixedtype); (2) learning the tabular data distribution directly using variational flow matching; (3) a systematic comparison of interpolation schemes; and (4) implementing stochastic dynamics in VFM.
At the application level, our work bridges flow matching with the operational needs of statistical agencies by providing a systematic analysis of the utility and risk of the synthetic data under different experimental conditions, recognising that algorithms optimising fidelity may implicitly amplify disclosure risk. The results demonstrate that flow matching can generate high-quality synthetic tabular data while protecting against statistical disclosure, thus offering a new option for responsible data generation.
2 Background
2.1 Flow Matching
Flow matching (FM) (lipman2023flowmatchinggenerativemodeling; liu2022flowstraightfastlearning; albergo2023stochasticinterpolantsunifyingframework) is a simulation-free generative modelling framework that learns a velocity field $v_{t}^{\theta}$ parameterised by $\theta\in\mathbb{R}^{p}$. This field defines a continuous transformation of a base distribution $p_{0}$ (typically a standard Gaussian) into a target distribution $p_{1}$ (empirical data) over a pseudo-time interval $t\in[0,1]$ through its induced ordinary differential equation.
Rather than directly estimating the target velocity field $u_{t}(x)$ – which we do not have access to – FM defines conditional distributions $p_{t}(x_{t}\mid x_{1})$ that make an assumption on the dynamics towards a fixed endpoint $x_{1}$, for which the conditional velocity is simple to compute. Under this formulation, the marginal velocity field is given by:
$u_{t}(x_{t})=\int u_{t}(x_{t}\mid x_{1})\frac{p_{t}(x_{t}\mid x_{1})p_{\text{data}}(x_{1})}{p_{t}(x_{t})}\,\mathrm{d}x_{1},$ (2.1)
where $u_{t}(x_{t}\mid x_{1})$ is the conditional velocity field. To avoid the computational cost of evaluating the integral in equation 2.1, FM learns the target velocity field through a Monte Carlo estimate of the conditional velocity, making the problem tractable. This approach is known as Conditional Flow Matching (CFM):
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t\sim[0,1],x_{1}\sim p_{\text{data}},x_{t}\sim p_{t}(x_{t}\mid x_{1})}\left[\left\lVert v_{t}^{\theta}(x_{t})-u_{t}(x_{t}\mid x_{1})\right\rVert_{2}^{2}\right].$ (2.2)
A common way to define this conditional distribution is through interpolation, i.e. define $x_{t}$ as a linear combination of $x_{1}$ and a noise sample $x_{0}$:
$x_{t}=\alpha_{t}x_{1}+\sigma_{t}x_{0}\sim p_{t}(x_{t}\mid x_{1})\quad\text{ for }\quad x_{0}\sim p_{0},\;x_{1}\sim p_{1},\;t\in[0,1].$ (2.3)
where $\alpha_{t}$, $\sigma_{t}$ are time-dependent coefficients that specify the trajectory. We can hence write down the corresponding conditional velocity:
$\begin{split}u(x_{t}\mid x_{1})&=\dot{\alpha}_{t}x_{1}+\dot{\sigma}_{t}\frac{x_{t}-\alpha_{t}x_{1}}{\sigma_{t}}=\underbrace{\Big(\dot{\alpha}_{t}-\frac{\dot{\sigma}_{t}}{\sigma_{t}}\alpha_{t}\Big)}_{A_{t}(x_{t})}x_{1}+\underbrace{\frac{\dot{\sigma}_{t}}{\sigma_{t}}}_{B_{t}(x_{t})}x_{t}.\end{split}$ (2.4)
Here, $A_{t}(x_{t})$ and $B_{t}(x_{t})$ are time-dependent coefficients that control the contributions of $x_{1}$ and $x_{t}$, respectively. This expression allows for efficient computation of the conditional velocity field without needing to perform numerical integration or simulation.
2.2 Variational Flow Matching
Although flow matching is already an efficient and simulation free generative modelling framework, it can struggle with multimodal data or data of heterogeneous types (zhai2025vfpvariationalflowmatchingpolicy). To address this limitation, eijkelboom2024variationalflowmatchinggraph put forward variational flow matching (VFM), which reinterprets flow matching using the perspective of variational inference.
Unlike standard FM - which is primarily designed for continuous data - VFM allows a more flexible learning mechanism that accommodates a wider range of data types, including tabular data (guzman-cordero2025exponential). VFM learns an approximate vector field via the posterior distribution of the data using a variational distribution $q^{\theta}_{t}$:
$v_{t}^{\theta}(x_{t}):=\int u_{t}(x_{t}\mid x_{1})q_{t}^{\theta}(x_{1}\mid x_{t})\,\mathrm{d}x_{1}.$ (2.5)
This formulation allows the model to flexibly approximate the true conditional velocity while allowing the incorporation of data-specific constraints (e.g. discrete data) through the choice of posterior. The training objective in VFM is to minimise the divergence between the true joint distribution and the variational joint distribution, measured in the Kullback-Leibler divergence (KLD):
$\mathcal{L}_{\text{MF-VFM}}(\theta)$ $=\mathbb{E}_{t}\left[\text{KL}\left(p_{t}(x_{t})p_{t}(x_{1}\mid x_{t})\,\Big\|\,p_{t}(x_{t})q_{t}^{\theta}(x_{1}\mid x_{t})\right)\right]$ (2.6)
$=-\mathbb{E}_{t,x_{1}x_{t}}\left[\log q_{t}^{\theta}(x_{1}\mid x_{t})\right]+\text{const}$ (2.7)
$=-\mathbb{E}_{t,x_{1},x_{t}}\left[\sum_{d=1}^{D}\log q_{t}^{\theta}(x_{1}^{d}\mid x_{t})\right]+\text{const},$ (2.8)
where the final line assumes a mean-field approximation (which does not restrict expressivity as posited in the theoretical analysis of eijkelboom2024variationalflowmatchinggraph) factorising the variational posterior across dimensions. Having learned the posterior distribution, the vector field can be computed using equation 2.9:
$v_{t}^{\theta}(x_{t})=\mathbb{E}_{q_{t}^{\theta}(x_{1}\mid x_{t})}\left[u_{t}(x_{t}|x_{1})\right]=\left(\dot{\alpha}_{t}-\frac{\dot{\sigma}_{t}}{\sigma_{t}}\alpha_{t}\right)\theta_{t}(x_{t})+\frac{\dot{\sigma}_{t}}{\sigma_{t}}x_{t}\\$ (2.9)
where $\theta_{t}(x_{t})$ is the mean of the variational distribution $q_{t}^{\theta}$. Crucially, this formulation recovers standard flow matching in the Gaussian case.
Stochastic Dynamics.
Beyond the deterministic ODE formulation, VFM also admits a stochastic counterpart that preserves the same marginal distribution $p_{t}(x)$ eijkelboom2024variationalflowmatchinggraph:
$\mathrm{d}x_{t}\;=\;\Big[u_{t}(x_{t})+\tfrac{g_{t}^{2}}{2}\,\nabla_{x_{t}}\log p_{t}(x_{t})\Big]\mathrm{d}t\;+\;g_{t}\,\mathrm{d}w_{t},$ (2.10)
where $g_{t}\!\geq\!0$ is a diffusion scheduler and $w_{t}$ is a standard Wiener process. The scheduler controls the stochasticity of the trajectory, but not the marginal path; Setting $g_{t}\!\equiv\!0$ recovers the ODE. We approximate this SDE using the variational distribution $q_{t}^{\theta}(x_{1}\mid x)$ as follows:
$\mathrm{d}x_{t}=\Big[v_{t}^{\theta}(x_{t})+\tfrac{g_{t}^{2}}{2}\,s_{t}^{\theta}(x_{t})\Big]\mathrm{d}t+g_{t}\,\mathrm{d}w_{t},$ (2.11)
where the drift and score are estimated as:
$v_{t}^{\theta}(x_{t})=\mathbb{E}_{q_{t}^{\theta}(x_{1}\mid x_{t})}[u_{t}(x_{t}\mid x_{1})],\qquad s_{t}^{\theta}(x_{t})=\mathbb{E}_{q_{t}^{\theta}(x_{1}\mid x_{t})}[\nabla_{x_{t}}\log p_{t}(x_{t}\mid x_{1})].$ (2.12)
This formulation supports approximate stochastic flow models, enabling richer models while still leveraging the variational structure for tractable learning.
2.3 Trajectory Setting in Flow Matching
There are two commonly used trajectories to define the probability path in flow matching (lipman2023flowmatchinggenerativemodeling).
The first is the optimal transport (OT) trajectory (liu2022flowstraightfastlearning; tong2024improvinggeneralizingflowbasedgenerative), where the mean and standard deviation of the conditional distribution evolve linearly over time. This produces a deterministic vector field with a fixed direction, resulting in straight-line (conditional) paths connecting the noise and data distributions (liu2022flowstraightfastlearning). OT-based trajectories are computationally straightforward, efficient, and easy to parameterise due to their time-invariant structure, as shown in equation 2.13
$\begin{split}x_{t}&=tx_{1}+(1-(1-\sigma_{\text{min}})t)x_{0}\\ u_{t}(x_{t}\mid x_{1})&=\frac{x_{1}-(1-\sigma_{\text{min}})x_{t}}{1-(1-\sigma_{\text{min}})t},\end{split}$ (2.13)
where $\sigma_{\text{min}}$ is added to ensure the target density has non-zero measure everywhere (or: we can imagine first convolving the data distribution with a small Gaussian).
The second type of trajectory is the variance-preserving (VP) path, derived from stochastic differential equations commonly used in diffusion models (lipman2023flowmatchinggenerativemodeling). In this approach, noise is gradually added to the data during the forward process, while the reverse process learns to de-noise using time-dependent vector fields. Unlike OT paths, VP diffusion trajectories evolve smoothly in both magnitude and direction over time. The parameters $\alpha_{t}$ and $\sigma_{t}$ are determined by cumulative functions of the noise scale function $\beta(t)=\beta_{\text{min}}+t(\beta_{\text{max}}-\beta_{\text{min}})$, as given in equation 2.14.
$\begin{split}\alpha_{t}&=e^{-\frac{1}{2}T(t)},\quad T(t)=\int_{0}^{1-t}\beta(s)ds,\quad\sigma_{t}=\sqrt{1-\alpha_{t}^{2}}\\ u_{t}(x_{t}|x_{1})&=-\frac{\dot{T}(t)}{2}\left[\frac{e^{-T(t)}x_{t}-e^{-\frac{1}{2}T(t)}x_{1}}{1-e^{-T(t)}}\right]\end{split}$ (2.14)
The choice of trajectory – or interpolation scheme – significantly affects data synthesis performance, which we explore in this paper. Depending on the data structure or modelling objectives, either trajectory can be more suitable. Additional trajectory formulations are provided in Appendix E.1.
2.4 TabSyn: Latent Diffusion for Tabular Data
TabSyn (zhang2024mixedtype) represents a state-of-the-art generative approach for mixed-type tabular data. It integrates a transformer-based Variational Autoencoder (VAE) with a score-based diffusion model operating in latent space. The design of TabSyn addresses several key challenges in tabular data synthesis including heterogeneous feature types, complex inter-feature dependencies, and the inefficiency of diffusion processes when employed on high-dimensional structured data.
The TabSyn framework comprises two stages. First, a VAE maps the raw tabular data into a continuous latent space. Second, a diffusion model is trained to capture the underlying distribution with this latent space. To handle different data types, TabSyn uses column-wise tokenisers and detokenisers, transformer-based encoders/decoders, and an adaptive $\beta$-VAE objective. The encoder learns structured latent representations, while the decoder ensures accurate and type-consistent data reconstruction.
Within the latent space, noise is introduced through a forward process defined as: $z_{t}=z_{0}+\sigma(t)\varepsilon$, with $\sigma(t)=t$ under a linear noise schedule. The corresponding reverse process is an SDE:
$\mathrm{d}z_{t}=-2\sigma_{t}\dot{\sigma}_{t}\nabla_{z}\log p_{t}(z_{t})\,\mathrm{d}t+\sqrt{2\sigma_{t}\dot{\sigma}(t)}\,\mathrm{d}\omega_{t}.$ (2.15)
where $\omega_{t}$ denotes a Wiener process. The model is trained using denoising score matching, with a neural network $\varepsilon_{\theta}(z_{t},t)$ that estimates the scoring function $\nabla_{z_{t}}\log p_{t}(z_{t})=-\varepsilon_{\theta}(z_{t},t)/\sigma(t)$.
A study by  zhang2024mixedtype showed that TabSyn surpasses traditional GAN and VAE based models as well as the latest diffusion models like TabDDPM and STaSy, across a variety of datasets and evaluation metrics - including unvariate distribution and pairwise correlation preservation, and downstream ML performance. Furthermore, its latent-space formulation allows for missing value imputation via latent inpainting, achieving competitive results in terms of data quality, diversity, and privacy preservation.
3 Designing Flow Matching for Tabular Data Synthesis
3.1 Motivation and Design Axes
Flow Matching provides a general recipe for learning deterministic or stochastic transformations between probability distributions. Yet, its effectiveness in practice depends not on a single equation, but on a set of design decisions that determine how distributions are represented, interpolated, and evolved over time. When applied to tabular data synthesis, these decisions interact in particularly complex ways: mixed continuous-categorical features challenge smooth trajectories, privacy-sensitive settings impose computational limits, and statistical agencies often require explainable and reproducible models. This section therefore reformulates Flow Matching as a design space rather than a single algorithmic recipe, and places tabular synthesis at its centre.
Our aim is to identify how four core axes – representation, learning target, trajectory, and dynamics – jointly shape the behaviour of tabular flow models.
  1. 1.
    Representation decides whether learning occurs in a continuous latent space that simplifies mixed data, or directly in the data space, preserving semantics.
  2. 2.
    Learning target distinguishes Conditional Flow Matching (velocity regression) from Variational Flow Matching (posterior regression), reflecting different ways of estimating the transport field.
  3. 3.
    Trajectory defines the interpolant between source and data distributions – commonly Optimal Transport (OT) or Variance Preserving (VP) paths – which control how signal and noise evolve over pseudo-time.
  4. 4.
    Dynamics determine whether the learned flow is integrated deterministically as an ODE or stochastically as an SDE, offering a continuum between precise transport and exploratory regularisation.
By systematically combining these axes, we focus on two complementary formulations. First, we introduce TabSynFlow, which learns deterministic ODE flows in latent space, substituting the diffusion process in TabSyn with a learned velocity field. We also evaluate TabbyFlow developed by guzman-cordero2025exponential which extends Variational Flow Matching directly to data space, supporting both numerical and categorical variables through hybrid Gaussian-categorical posteriors.
We consider a dataset of mixed-type tabular observations $x=(x_{\text{num}},x_{\text{cat}}),$ where $x_{\text{num}}\in\mathbb{R}^{D_{\text{num}}}$ are continuous features and $x_{\text{cat}}\in\{1,\dots,K_{d}\}^{D_{\text{cat}}}$ are categorical variables. The empirical data distribution is denoted by $p_{1}(x)$, while a simple prior $p_{0}(x)$, typically an isotropic Gaussian, serves as the source distribution. The goal of flow matching is to learn a continuous transformation that maps $p_{0}$ to $p_{1}$ through a parameterised time-dependent velocity field $v_{t}^{\theta}(x_{t})$ defined on the interpolation in equation 2.3. The coefficients $(\alpha_{t},\sigma_{t})$ specify the probability path (for example, Optimal Transport or Variance Preserving trajectories) and induce a family of intermediate marginals $p_{t}(x_{t})$. The generative process is recovered by integrating either a deterministic ordinary differential equation (ODE) or a stochastic differential equation (SDE) via equation 2.10. The two instantiations below, TabbyFlow and TabSynFlow, correspond to different modelling choices for the representation space and the learning target.
3.2 TabbyFlow: Variational Flow Matching in Data Spaces
TabbyFlow (guzman-cordero2025exponential) extends the Variational Flow Matching (VFM) framework of eijkelboom2024variationalflowmatchinggraph to structured tabular data containing both numerical and categorical features. This formulation enables direct modelling in the data space, avoiding the need for an intermediate latent representation, while explicitly accounting for heterogeneous variable types within the flow dynamics.
The TabbyFlow objective treats each feature type according to its statistical nature: continuous variables are modelled using a Gaussian distribution, while categorical variables are modelled using categorical distributions. Let $x=(x_{\text{num}},x_{\text{cat}})$ denote a sample comprising $D_{\text{num}}$ numerical and $D_{\text{cat}}$ categorical features. Each categorical feature is one-hot encoded, yielding total dimensionality
$D=D_{\text{num}}+\sum_{d=1}^{D_{\text{cat}}}K_{d},$ (3.1)
where $K_{d}$ is the number of categories for the $d$-th categorical variable.
Numerical Variables.
Following Theorem 3 of eijkelboom2024variationalflowmatchinggraph, we assume a mean-field variational posterior $q_{t}^{\theta}(x_{\text{num},1}\mid x_{t})$ with Gaussian structure parameterised by mean $\theta_{t}(x_{t})$ and variance $0.5A_{t}(x_{t})^{-2}$. The VFM loss for numerical variables is
$\mathcal{L}_{\text{VFM-num}}(\theta)=\mathbb{E}_{t,x_{t},x_{1}}\left[\|A_{t}(x_{t})(x_{\text{num},1}-\theta_{t}(x_{t}))\|_{2}^{2}\right].$ (3.2)
However, we found that using $0.5A_{t}(x_{t})^{-2}$ as variance yielded lower performance in our experiment. To mitigate this, we implemented a relaxed variance of $0.5A_{t}(x_{t})^{-1}$. This adjustment slightly inflates the variance, therefore acting as a mild regulariser. The empirical results can be seen in appendix D.1.
Categorical Variables.
For categorical features, TabbyFlow models the variational posterior as a product of categorical distributions. Let $\theta_{t}^{dk}(x_{t})$ denote the predicted probability of category $k$ in the $d$-th feature, obtained via softmax in the network output layer. The loss minimises cross-entropy with the one-hot targets $x_{\text{cat},1}^{d}$:
$\mathcal{L}_{\text{VFM-cat}}(\theta)=-\mathbb{E}_{t,x_{t},x_{1}}\left[\sum_{d=1}^{D_{\text{cat}}}\sum_{k=1}^{K_{d}}\mathbb{I}[x_{\text{cat},1}^{d}=k]\log\theta_{t}^{dk}(x_{t})\right].$ (3.3)
Unified Objective.
Combining both loss terms yields the overall TabbyFlow objective:
$\begin{split}\mathcal{L}_{\text{TabbyFlow}}(\theta)=&\;\mathbb{E}_{t,x_{t},x_{1}}\left[\|\sqrt{A_{t}(x_{t})}(x_{\text{num},1}-\theta_{t}(x_{t}))\|_{2}^{2}\right]\\ &\;-\mathbb{E}_{t,x_{t},x_{1}}\left[\sum_{d=1}^{D_{\text{cat}}}\sum_{k=1}^{K_{d}}\mathbb{I}[x_{\text{cat},1}^{d}=k]\log\theta_{t}^{dk}(x_{t})\right].\end{split}$ (3.4)
Sampling and Dynamics.
During sampling, a noise sample $\tilde{x}_{0}$ is drawn from $p_{0}$, and the trained network predicts $\theta_{t}(x_{t})$ at each time step. Because TabbyFlow does not directly output a velocity field, it is recovered via
$v_{t}^{\theta}(x_{t})=\left(\dot{\alpha}_{t}-\frac{\dot{\sigma}_{t}}{\sigma_{t}}\alpha_{t}\right)\theta_{t}(x_{t})+\frac{\dot{\sigma}_{t}}{\sigma_{t}}x_{t},$ (3.5)
and, for the stochastic variant, the score function
$s_{t}^{\theta}(x_{t})=-\frac{x_{t}-\alpha_{t}\theta_{t}(x_{t})}{\sigma_{t}^{2}}.$ (3.6)
These components define the ODE or SDE dynamics used for synthesis (Algorithms 34 in Appendix B.2). TabbyFlow thus enables a unified treatment of numerical and categorical features within a single variational flow, generalising continuous-time generative modelling to structured tabular domains.
3.3 TabSynFlow: Flow Matching in Latent Spaces
TabSynFlow extends latent-space generative modelling by replacing the diffusion component in TabSyn with a deterministic probability flow trained via Conditional Flow Matching (CFM). The model learns a time-dependent velocity field that transports a simple prior to the target latent distribution through an ODE, enabling a direct and continuous transformation during both training and sampling. This deterministic integration stabilises training and accelerates synthesis relative to iterative stochastic denoising, while retaining high-fidelity latent reconstructions through the encoder–decoder pathway.
TabSynFlow follows a two-stage framework. In the first stage, a VAE maps the raw tabular data, comprising both continuous and categorical features, into a continuous latent space. Subsequently, flow matching is applied to model the latent distribution $p(z_{1})$, replacing the score-based diffusion model used in TabSyn. Gaussian noise samples $z_{0}\sim\mathcal{N}(0,I)$ are continuously transformed into latent variables $z_{1}$ through an ODE parameterised by $v_{t}^{\theta}(z_{t})$. The conditional velocity $u_{t}(z_{t}\mid z_{1})$ is computed using Eq. equation 2.4, and the loss becomes
$\mathcal{L}_{\text{TabSynFlow}}(\theta)=\mathbb{E}_{t,z_{1},z_{t}}\left[\|v_{t}^{\theta}(z_{t})-u_{t}(z_{t}\mid z_{1})\|_{2}^{2}\right].$ (3.7)
After training, sampling proceeds by drawing $z_{0}\sim\mathcal{N}(0,I)$ and integrating the learned ODE until $t=1$. The resulting latent representation $z_{1}$ is then decoded to obtain a synthetic record $\tilde{x}$, followed by inverse transformation to the original feature space. Algorithms 12 in Appendix B.1 describe the procedure in detail.
Operating in latent space grants TabSynFlow efficient and stable training while retaining the expressive power of flow-based models. It supports fast, deterministic sampling and interpretable trajectories, making it a computationally efficient alternative to diffusion-based approaches for tabular data synthesis.
4 Experimental Setting
Table 1 summarises the datasets used in our empirical evaluation. Following  ran2024multiobjective and  Little2024synthetic, we used four census datasets (from the UK, Fiji, Canada and Rwanda), plus one additional census data from Indonesia. We also added two further databases from UCI which are standard benchmarks for tabular data kotelnikov2023tabddpm. Unlike ran2024multiobjective, we treat age as continuous to exploit its natural ordering. Also, this choice remains compatible with TabDDPM/TabSyn, which operate on mixed-type data. To ensure a consistent and fair comparison, we adopt the same convention for TabSynFlow and TabbyFlow.
All the models in this study use a multilayer perceptron (MLP) architecture comprising four hidden layers with sizes [1024, 2048, 2048, 1024], configuration used by  zhang2024mixedtype. TabSyn and TabSynFlow employ Transformer-based architectures in their VAE counterparts, diverging from the standard MLP-based tabular VAE design (xu2019modeling).
Training was conducted with a batch size to 4096, and the ODE was integrated until $t=1$ using 100 steps  [^1] Unless otherwise specified, this is the default setup. Models were trained for up to 10,000 epochs, with early stopping based on the lowest observed training loss, which is similar to previous studies (zhang2024mixedtype; guzman-cordero2025exponential), although those studies capped training at 8000 epochs. The integration for synthesis uses Euler method. For sampling from the SDE, we set $g_{t}=\sigma_{t}$.
We focused the synthetic tabular data evaluation on preserving the statistical properties of the original data (utility) while minimising the potential to disclose sensitive information about individuals (disclosure risk) (taub2019the; elliot2023samples; Little2022comparing). Utility values range from 0 (no preservation of statistical properties) to 1 (perfect preservation), measuring how well synthetic data supports the same analyses as the original. Disclosure risk ranges from 0 (no added risk beyond baseline) to 1 (maximum disclosure risk), quantifying the probability that an adversary can correctly match sensitive information using synthetic data. Detailed evaluation methods are provided in Appendix C.
Table 1: Datasets used for the empirical study.
Abbr. Dataset Name #Obs. #Num. #Cat.
Variables Variables
UK UK Census 104267 1 14
CA Canada Census 32149 4 21
FI Fiji Census 84323 1 18
RW Rwanda Census 31455 1 12
ID Indonesia Census 177429 1 12
AD Adult 48842 5 10
CH Churn Modelling 10000 4 7
5 Results
In this section, we present and analyse the empirical findings of our study, highlighting the comparative performance of diffusion and flow-matching approaches for tabular data synthesis. As seen from the objective function, all algorithms optimise for utility; disclosure risk is measured post-hoc. We begin by establishing TabDDPM and TabSyn as strong baselines (Section 5.1). We then examine the number of function evaluations (NFEs), evaluating the utility and risk dynamics between computational efficiency (Section 5.2). Next, we investigate the impact of integration time ($t_{ode}$) on model performance, with specific attention to the evolution of probability paths (Section 5.3). Finally, we evaluate the effect of incorporating stochastic sampling within the VFM framework (Section 5.4). Collectively, these analyses provide a picture of how modelling choices shape the quality, efficiency, and reliability of synthetic tabular data generation.
5.1 Main Comparison
In this section, we present results for the census datasets which have been used for benchmarking in previous studies  (e.g. ran2024multiobjective; Little2024synthetic). Table 2 summarises the comparative performance of six synthetic data generation algorithms across UK, CA, FI, and RW datasets.
TabDDPM exhibits highly inconsistent results, excelling in UK while substantially underperforming in CA, FI, and RW. On the other hand, TabSyn demonstrates robust utility across all datasets, achieving values above $0.6$ in RW and exceeding $0.7$ in UK, CA, and FI, making it a reliable choice for utility-driven synthesis.
However, TabbyFlow proves to be the most effective model for these benchmarks. The OT trajectory delivers the highest utility in UK and CA, while the VP trajectory attains the best scores in FI and RW. Notably, TabbyFlow-VP consistently emerges as the second-best in utility, and further distinguishes itself by achieving lower risk, particularly in FI and RW.
From regular flow matching side, TabSynFlow-OT outperforms TabSyn in UK and RW, attaining lower disclosure risk in their high utility synthetic data, indicative of improved synthesis quality. In CA and FI, TabSynFlow demonstrates a classic risk–utility trade off. Moreover, TabSynFlow-VP achieves the most favorable risk profiles in CA, FI, and RW at the cost of lower utility.
In summary, TabbyFlow—with both OT and VP trajectories—consistently attains the highest overall average scores — calculated via average of utility and 1-risk — on the main census benchmark, achieving values of 0.7038 (OT) and 0.7031 (VP). TabSyn and TabSynFlow-OT also perform competitively, with average scores of 0.6742 and 0.6716, respectively, ranking just below the TabbyFlow models. In comparison, TabSynFlow-VP yields a lower average score of 0.6150, while TabDDPM lags behind at 0.5083. These findings underscore the robustness of flow matching algorithms, particularly the TabbyFlow family, in effectively high data utility and less disclosure risk. The findings also highlight the critical role of trajectory choice (OT or VP) that modulates the relationship in some datasets, particularly the risk as post-hoc consequence. Results for the additional datasets (Indonesia, Adult, and Churn) are available in Appendix D, and further analyses of alternative interpolants for both TabSynFlow and TabbyFlow are provided in Appendix E.2 and Appendix E.3, respectively.
Table 2: Comparative Performance of six synthetic data generation algorithms across four common census datasets (UK, CA, FI, RW). Bold and underline indicate the first and second best performing algorithms for each dataset, respectively. The table emphasises best performance of TabbyFlow in common census benchmarks.
Evaluation Algorithm UK CA FI RW
[…]
5.2 Number of Function Evaluations between Tabsyn and Flow Matching Models
The analysis of computational efficiency presented in Figure 1 illustrates the relationship between the number of function evaluations (NFEs) and the performance of synthetic data generation models, highlighting the relationship between computational cost and data quality.
From the figure, it can be seen that the FM models generally converge after approximately 100 NFEs, indicating that using low steps in FM is sufficient. On the other hand, TabSyn requires more than 100 steps to outperform FM models in terms of utility, indicating a higher computational burden, in addition to costly training of VAE+diffusion.
In low-NFE settings, where flow-based models are designed to operate efficiently, TabSyn performed poorly. TabbyFlow-VP whilst yielding lower utility also achieves lower disclosure risk, and still outperforms TabSyn when NFEs are below 64. This early-stage advantage positions flow matching models, particularly using the OT path as promising candidates for applications operating under strict computational constraints.
In summary, while TabSyn can achieve competitive utility, it does so at the cost of significantly higher NFEs. Flow-matching models offer a more efficient alternative, especially in scenarios where computational efficiency is critical.
Figure 1: Average utility ($\uparrow$) and disclosure risk ($\downarrow$) as a function of the number of function evaluations (NFEs) for TabSyn and flow matching models (TabSynFlow-OT, TabSynFlow-VP, TabbyFlow-OT, TabbyFlow-VP) averaged across four datasets Solid lines represent utility, while dotted lines represent risk. Flow-matching models converge after approximately 100 NFEs, achieving competitive utility at substantially lower computational cost compared to TabSyn.
5.3 Flow Matching Results based on Integration Time
In standard flow matching implementations, the ODE is fully integrated; however, early termination is possible in practice. Figure 2 provides an analysis of the effect of ODE integration times ($t_{\text{ode}}$) on utility and risk, averaged across all datasets. For TabSynFlow-OT, the utility remains consistently high throughout integration times, demonstrating robust performance even with early stopping. As in $t_{\text{ode}}\rightarrow 1.0$, the performance across all methods tends to plateau.
In contrast, both TabSynFlow and TabbyFlow using the VP path exhibit a lower initial utility at $t_{\text{ode}}=0.6$, suggesting that VP introduces disruptive noise early in the process. Utility and risk increase substantially when $t_{\text{ode}}>=0.6$, indicating that VP require full integration to achieve competitive utility. From these results, we can see that the consistent superiority of OT paths across integration times highlighting their robustness where early termination is desirable for computationally constrained applications. Dataset-specific results provided in Appendix D.4 show deviations around the average, particularly for $t_{\text{ode}}=[0.9,1]$, emphasising the importance of considering dataset characteristics in selecting integration parameters.
Figure 2: Average utility ($\uparrow$) and risk ($\downarrow$) against ODE integration time ($t_{\text{ode}}$) for TabSynFlow and TabbyFlow using OT and VP paths. Colours and line styles indicate method and path (VP light and solid, OT dark and dashed). OT paths enable early stopping without significant utility loss, whereas VP paths require full integration to achieve competitive performance.
5.4 SDE Implementation of Variational Flow Matching
In this subsection, we investigate whether stochastic differential equation (SDE) sampling yields results comparable to ODE sampling. Tables 4 and 4 present a comparative analysis of ODE and SDE formulations, with additional results provided in Appendix E.4 for alternative diffusion coefficients.
In general, it can be seen that the SDE approach performs as expected theoretically, producing marginal distributions similar to those generated by ODE  eijkelboom2024variationalflowmatchinggraph while, in some cases, producing superior performance. This is evidenced by the close alignment of utility and risk values across both formulations. This empirical validation confirms the mathematical foundation presented in the paper.
Beyond theoretical equivalence, SDE sampling offers a practical advantage by improving the ODE results in some cases. For example, in the Fiji and Rwanda datasets across both experimental setups, SDE-based VFM simultaneously produced high utility synthetic data with lower risk - a dual improvement that is rare in practice, where gains typically only occur in one metric.
Table 3: Comparison of average utility and risk between ODE and SDE performance for TabbyFlow-OT.
Dataset ODE SDE
Utility Risk Utility Risk
UK 0.8333 0.4889 0.8306 0.4916
Indonesia 0.6671 0.4114 0.6629 0.4081
Canada 0.7718 0.3207 0.7785 0.3152
Fiji 0.7263 0.5705 0.7394 0.5664
Rwanda 0.7284 0.5379 0.7353 0.5361
Adult 0.7720 0.5443 0.7731 0.5508
Churn 0.7966 0.0773 0.7986 0.0784
Table 4: Comparison of average utility and risk between ODE and SDE performance for TabbyFlow-VP.
Dataset ODE SDE
Utility Risk Utility Risk
UK 0.8066 0.4765 0.7903 0.4543
Indonesia 0.6488 0.3918 0.6386 0.5005
Canada 0.7472 0.3008 0.7414 0.2926
Fiji 0.7451 0.5588 0.7859 0.5343
Rwanda 0.7358 0.5180 0.7430 0.5032
Adult 0.7581 0.5677 0.7514 0.5039
Churn 0.8067 0.0753 0.7941 0.1079
6 Discussion and Concluding Remarks
6.1 Discussion
The experimental results demonstrate the performance characteristics of flow-based deep generative models for tabular data synthesis, highlighting how utility and disclosure risk vary across algorithms and datasets.
Performance Comparison (Section 5.1). Flow matching models consistently outperform TabDDPM on most datasets, with TabbyFlow generating synthetic data of the highest utility in nearly all cases. This increased fidelity is generally accompanied by elevated disclosure risk, reflecting a positive empirical correlation between utility and risk. Our most interesting finding is while it is not commonly used in FM studies, using VP path in TabbyFlow has potential to produce better optimised synthetic data with higher utility but lower disclosure risk. TabSyn remains a reliable benchmark, while TabSynFlow-OT shows targeted improvements over TabSyn in selected cases but not universally. These results highlight flow matching—particularly TabbyFlow—as a viable alternative for tabular data synthesis, with trajectory choice gives different performance in terms of data utility and disclosure risk.
Efficiency via NFEs (Section 5.2). Flow matching models, particularly TabbyFlow-OT and TabSynFlow-OT, demonstrate strong utility with a relatively small number of function evaluations (NFEs), typically between 64 and 128, whereas TabSyn requires substantially more NFEs to reach comparable performance. Notably, TabbyFlow can achieve near-optimal utility by NFEs 128, and even at very low budgets (four evaluation steps), meaningful results are observed. The actual impact of NFEs, however, depends heavily on data size, feature complexity, and model architecture. For instance, generating small datasets like CH is fast, but larger datasets such as UK or ID require more time and batch-wise processing to avoid out-of-memory errors, particularly on personal computers. Overall, lower NFEs values generally reflect greater computational efficiency, making flow-based models especially attractive for resource-constrained environments.
Integration Time and Early Stopping (Section 5.3). OT paths outperform VP paths in both TabSynFlow and TabbyFlow, maintaining high utility earlier in the generation process. VP paths require full integration to achieve competitive performance. However, both trajectories are prone to reduced utility and / or increased risk when the integration time approaches one ($t_{ode}\to 1$). Practical strategies to mitigate this include early stopping ($t_{ode}\leq 0.9$) or velocity clipping222e.g., clip(1-1/(1-t), 0, m), m is the maximum value allowed. In general, OT trajectories provide a more stable and reliable approach to tabular synthesis.
Deterministic vs. stochastic sampling (Section 5.4). The application of SDE solvers within the VFM framework presents a more complex picture than initially hypothesised. While it is theoretically established that SDE recovers the same marginal distributions as ODE, our empirical results reveal numerical differences. In particular, on datasets such as Fiji and Rawnda, SDE sampling led to a simultaneous increase in utility and reduction in risk. This dual improvement suggests that the controlled stochasticity of the SDE can act as a beneficial regulariser or enable more effective exploration of the data space, rather than merely introducing destructive noise. However, since performance gains were not observed across all datasets, the utility of SDEs appears to be data-dependent. Thus, the choice between deterministic and stochastic samplers is not about universal dominance, but about suitability to specific dataset.
6.2 Concluding Remarks
This study has presented a comprehensive evaluation of flow matching models for tabular data synthesis, comparing them with state-of-the-art diffusion-based approaches. Our results demonstrate that flow matching—particularly TabbyFlow—consistently achieves higher utility than diffusion baselines, while disclosure risk follows patterns determined by the empirical relationship between model fidelity and privacy exposure, rather than optimisation for risk reduction. Across diverse datasets, the OT trajectory consistently achieves balanced performance with high utility and reduced sensitivity to integration time, while the VP trajectory enhances privacy protection with lower risk at the expense of some utility. In practice, trajectory selection should be guided by specific application requirements.
TabbyFlow further distinguishes itself through its computational efficiency, achieving competitive results with less function evaluations (NFEs). In practice, we recommend the OT path as the default configuration, with early stopping ($0.9\leq t_{\text{ode}}<1.0$) and 64–128 NFEs to balance between synthetic data quality and computational cost. Even ultra-low settings (4-8 NFEs) remain viable for resource-constrained environments.
The introduction of stochasticity via SDE solvers produced nuanced outcomes. While SDE formulations theoretically preserve marginal distributions, empirical performance varied between datasets. In several cases, SDE sampling simultaneously improved utility and reduced risk, demonstrating that stochastic dynamics can act as a beneficial regulariser or enhance exploration of the data space. However, these benefits were not universal, indicating that the effectiveness of SDE is context-dependent. Therefore, SDE-based sampling should be considered an optional empirical extension, subject to dataset-specific validation.
Overall, flow matching with the OT paths emerges as an efficient approach to tabular data synthesis. Future work could explore hybrid trajectories, incorporate formal privacy guarantees such as differential privacy, and investigate more thoroughly the dataset-specific factors that drive the effectiveness of both ODE and SDE solvers.
Acknowledgments
B.I. Nasution is supported by the Indonesia Endowment Fund for Education Agency (LPDP) scholarship.
Appendix A Related Work
Diffusion Models for Tabular Data. Denoising diffusion probabilistic models (DDPMs) have shown great success across domains like vision and language, but their application to tabular data poses unique challenges due to feature heterogeneity and small dataset sizes. TabDDPM (kotelnikov2023tabddpm) addresses this through a dual process: Gaussian diffusion for continuous variables and multinomial diffusion for categorical features. It achieves state-of-the-art (SOTA) utility and privacy performance, outperforming the GAN and VAE baselines. However, its computational cost remains high because of iterative sampling and separate handling of feature types, which may weaken correlation modelling. TabSyn (zhang2024mixedtype) employs a transformer-VAE architecture followed by a diffusion model, while the continuous diffusion model for mixed type tabular data (CDTD) (mueller2025continuous) is a score-matching diffusion model that noises continuous values and categorical embeddings and samples via ODE integration. These two methods improve the utility of tabular data, particularly compared to TabDDPM and other recent SOTAs.
FM and its Variational Extensions. Flow matching (FM) (lipman2023flowmatchinggenerativemodeling; liu2022flowstraightfastlearning; albergo2023stochasticinterpolantsunifyingframework) offers an alternative to diffusion by learning the deterministic dynamics of probability paths. FM has also been developed in terms of latent models for image generation (dao2023flowmatchinglatentspace; schusterbauer2024boosting). In addition, VFM (eijkelboom2024variationalflowmatchinggraph) reformulates FM as a variational inference problem, allowing flow matching through a variational distribution. Consequently, VFM can be implemented to mixed-type data, such as tabular data, through mean-field variational inference. On the other hand, (guzman-cordero2025exponential) extended the VFM specific for the exponential family that also supports mixed-type variables through moment matching and implemented it in tabular data, called TabbyFlow. In addition to tabular, VFM has also been implemented for image generation maticsan2025purrception, controlled molecular generation (eijkelboom2025controlled), non-Euclidean geometry  (zaghen2025variationalflowmatchinggeneral; zaghen2025riemannianvariationalflowmatching), and climate modeling (finn2025generative).
FM Trajectory. Trajectory design is of paramount importance in the field of FM. The paths of optimal transport (OT) are widely preferred because of their straightforward linear interpolation. In contrast, the variance-preserving (VP) path, which takes inspiration from diffusion models, is somewhat underrated (lipman2023flowmatchinggenerativemodeling; liu2022flowstraightfastlearning). Recent studies also introduced different trajectories such as cosine albergo2023buildingnormalizingflowsstochastic, logit-normal labs2025flux1kontextflowmatching. Although the theoretical framework of FM accommodates both types of path, there is a notable gap in empirical evaluation concerning the practical application of different paths, particularly in the realm of tabular data. To address this gap, we present the potential of different paths for tabular data synthesis, despite the fact that OT is the primary implementation.
Appendix B Algorithms
B.1 TabSynFlow
Algorithm 1 An iteration of TabSynFlow training (flow matching part)
Input: Latent representation from VAE $\mathcal{Z}_{1}$, time function for mean $\alpha_{t}$, time function for variance $\sigma_{t}$, conditional velocity field $\mathrm{NN}(.)$
[…]
5:Calculate $u_{t}(z_{t}|z_{1})$ using equation 2.4
6:Predict $v_{t}^{\theta}(z_{t})=\mathrm{NN}(z_{t},t)$
7:Calculate $\mathcal{L}_{\mathrm{TabSynFlow}}$
Algorithm 2 TabSynFlow sample generation
Input: Trained conditional velocity field $\mathrm{NN}(.)$, discrete step interval $\Delta$, trained decoder from VAE $\mathrm{Dec}(.)$
[…]
2:$t=0$
3:while $t\leq 1$ do
4:  $\tilde{z}_{t+\Delta}=\tilde{z}_{t}+\Delta\cdot\mathrm{NN}(\tilde{z}_{t},t)$
5:  $t\textrm{+=}\Delta$
6:end while
[…]
Return $\tilde{\mathcal{D}}$
B.2 TabbyFlow
Algorithm 3 TabbyFlow training iteration
Input: Preprocessed data, Function of time for mean $\alpha_{t}$, function of time for variance $\sigma_{t}$, variational parameter model $\mathrm{NN}(.)$
[…]
5:Predict $\theta_{t}(x_{t})=\mathrm{NN}(x_{t},t)$
6:Calculate $\mathcal{L}_{\mathrm{TabbyFlow}}(\theta)$
Algorithm 4 TabbyFlow sample generation
Input: Time function for mean $\alpha_{t}$, time function for variance $\sigma_{t}$, variational parameter model $\mathrm{NN}(.)$, discrete step interval $\Delta$, noise schedule $g_{t}$ (for SDE)
[…]
2:$t=0$
3:while $t\leq 1$ do
4:  $\tilde{\theta}_{t}(\tilde{x}_{t})=\mathrm{NN}(\tilde{x}_{t},t)$
5:  Calculate $v_{t}^{\theta}(\tilde{x}_{t})$ using equation 3.5
6:ODE:
7:  $\tilde{x}_{t+\Delta}=\tilde{x}_{t}+\Delta\cdot v_{t}^{\theta}(\tilde{x}_{t})$
8:SDE:
9:  Calculate $s_{t}^{\theta}(x)$ using equation 3.6
10:  $\tilde{x}_{t+\Delta}=\tilde{x}_{t}+\Delta\Big[v_{t}^{\theta}(\tilde{x}_{t})+\tfrac{g_{t}^{2}}{2}\,s_{t}^{\theta}(\tilde{x}_{t})\Big]+g_{t}\sqrt{\Delta}\,\mathcal{N}(0,I)$
11:  $t\textrm{+=}\Delta$
12:end while
[…]
Return $\tilde{\mathcal{D}}$
Appendix C Evaluation Details
This section provides detailed information on the evaluation metrics and implementation procedures used to assess the quality and privacy aspects of the synthetic tabular data generated in our experiments. This additional context aims to ensuring transparency and reproducibility, allowing readers to fully understand how each score is derived and how our experimental protocol aligns with established practices in tabular data synthesis evaluation.
C.1 Utility Metrics
Utility measures the extent to which synthetic data faithfully reproduces the statistical relationships and analytical properties of the original data. We employ two complementary utility metrics:
C.1.1 Ratio of Counts (ROC)
ROC evaluates the fidelity of frequency tables and cross-tabulations by comparing cell counts between original and synthetic data. For each cell, the ROC is:
$\mathrm{ROC}_{\text{cell}}=\frac{\min(\mathcal{Y}_{count},\mathcal{Y}^{\prime}_{count})}{\max(\mathcal{Y}_{count},\mathcal{Y}^{\prime}_{count})},$ (C.1)
where $\mathcal{Y}_{count}$ and $\mathcal{Y}^{\prime}_{count}$ denote cell counts in the original and synthetic data respectively. A value of 1 indicates perfect agreement; 0 indicates extreme disagreement.
ROC is computed across two types of tabulations:
  • Univariate: Frequency distribution of each key variable independently
  • Bivariate: Cross-tabulation of all pairs of key variables
The final $\text{ROC}_{\text{uni}}$ score is the mean across all univariate tables, and $\text{ROC}_{\text{biv}}$ is the mean across all bivariate tables, i.e. cross-tabulation.
C.1.2 Confidence Interval Overlap (CIO)
CIO assesses how well statistical inference (specifically logistic regression) conducted on synthetic data recovers the confidence intervals from the original data. For each regression coefficient, CIO calculates the normalized overlap of 95% confidence intervals:
$\mathrm{CIO}_{k}=0.5\left[\frac{\mathrm{overlap}}{u-l}+\frac{\mathrm{overlap}}{u^{\prime}-l^{\prime}}\right],$ (C.2)
where $[l,u]$ and $l^{\prime},u^{\prime}]$ are the confidence intervals from original and synthetic models respectively, and overlap = $\min(u,u^{\prime})-\max(l,l^{\prime})$. A CIO of 1 indicates identical intervals, whereas CIO < 0 indicates no overlap, therefore it is truncated to zero (elliot2023samples). The final CIO score is the mean overlap across all regression coefficients and all fitted models.
C.1.3 Aggregate Utility Score
The final utility score for each dataset is computed as:
$\text{Utility}=\frac{\mathrm{ROC_{uni}}+\mathrm{ROC_{biv}}+\mathrm{CIO}}{3}$ (C.3)
For each dataset, we compute ROC across all univariate and bivariate frequency tables, and CIO for regression models predicting a defined target variable. Mainly, we used logistic regression, except for target variable CreditScore and EstimatedSalary in Churn data that uses linear regression since they are continuous variables. The pseudocode to compute ROC and CIO can be seen in Algorithms 5 and 6.
Algorithm 5 ROC calculation
Input: Original data $\mathcal{D}$, synthetic data $\mathcal{D}^{\prime}$, key variables $K$
1:for each variable $k$ in $K$ do
2:  for each category $c$ of $k$ do
3:   Compute $\text{ROC}_{\text{uni}}[c]=\frac{\min(\mathcal{Y}_{count}(c),\mathcal{Y}^{\prime}_{count}(c))}{\max(\mathcal{Y}_{count}(c),\mathcal{Y}^{\prime}_{count}(c))}$
4:  end for
5:end for
[…]
7:
8:for each pair $(k_{1},k_{2})$ in $K$ do
9:  for each cell $(c_{1},c_{2})$ do
10:   Compute $\text{ROC}_{\text{biv}}[c_{1},c_{2}]=\frac{\min(\mathcal{Y}_{count}(c_{1},c_{2}),\mathcal{Y}^{\prime}_{count}(c_{1},c_{2}))}{\max(\mathcal{Y}_{count}(c_{1},c_{2}),\mathcal{Y}^{\prime}_{count}(c_{1},c_{2}))}$
11:  end for
12:end for
[…]
14:Return $\text{ROC}_{\text{uni}},\;\text{ROC}_{\text{biv}}$
Algorithm 6 CIO calculation
Input: Original data $\mathcal{D}$, synthetic data $\mathcal{D}^{\prime}$, key variables $K$, target variable $T$
1:Fit regression model (predictors $K$, target $T$) on $\mathcal{D}$
2:Fit regression model (predictors $K$, target $T$) on $\mathcal{D}^{\prime}$
3:for each coefficient $i$ do
4:  Compute $[l,u]$ from model on $\mathcal{D}$
5:  Compute $l^{\prime},u^{\prime}]$ from model on $\mathcal{D}^{\prime}$
6:  $\text{overlap}=\max(0,\min(u,u^{\prime})-\max(l,l^{\prime}))$
7:  $\text{CIO}_{i}=0.5\left[\frac{\text{overlap}}{u-l}+\frac{\text{overlap}}{u^{\prime}-l^{\prime}}\right]$
8:end for
[…]
10:Return CIO
C.2 Disclosure Risk Metric
We measure disclosure risk as the probability that an adversary, possessing the quasi-identifier (key variables, $K$), can correctly guess the sensitive value (target $T$) for a record by matching between the synthetic and original data. This approach, adapted from taub2019the, quantifies the additional privacy risk introduced by synthetic data at a per-record granularity, adjusting for the baseline (based on a random draw from $T$), defined by the distribution of $T$ within each equivalence class (combination of $K$) in the original data.
For each synthetic record, we first identify the group of records in the synthetic data that share the same values for the key variables, as well as the more specific group that also matches on the target variable. The proportion is then calculated as the number of records sharing both key and target values divided by the number sharing only the key values. This proportion is used to filter which synthetic records are included in the TCAP calculation: only those in groups where the proportion meets or exceeds a specified threshold (represented by the parameter $\tau$ - which represents an assumed lower bound on the acceptable level of uncertainty for the adversary) are considered, while records in less concentrated groups are excluded. This filtering is applied solely within the synthetic data; the actual TCAP risk for each retained record is subsequently calculated based on the target distribution in the original data for the corresponding key group.
Targeted Correct Attribution Probability (TCAP) estimates the probability that the adversary can match the synthetic target $T^{\prime}_{j}$ using the equivalence class from the original data:
$TCAP^{\prime}_{j}=P_{\mathcal{D}}(T^{\prime}_{j}|K^{\prime}_{j})=\frac{\sum_{i=1}^{n}\mathbb{I}[T_{i}=T^{\prime}_{j},\,K_{i}=K^{\prime}_{j}]}{\sum_{i=1}^{n}\mathbb{I}[K_{i}=K^{\prime}_{j}]}$ (C.4)
Intuitively, $TCAP^{\prime}_{j}$ shows the chance that target $T^{\prime}_{j}$ could arise in the real data for key $K^{\prime}_{j}$—that is, it reflects the “true match” risk based on empirical frequencies in the original dataset.
Within-Equivalence Class Attribute Probability (WEAP) serves as a baseline: if the adversary could only guess randomly within the equivalence class $K^{\prime}_{j}$, their success probability would be
$WEAP_{j}=P_{\mathcal{D}}(T_{j}|K_{j})=\frac{\sum_{i=1}^{n}\mathbb{I}[T_{i}=T_{j},K_{i}=K_{j}]}{\sum_{i=1}^{n}\mathbb{I}[K_{i}=K_{j}]}$ (C.5)
This is identical in calculation to $TCAP^{\prime}_{j}$, but conceptually $WEAP_{j}$ captures the average success rate without access to synthetic information—i.e., a “random guess” adversary.
Per-record disclosure risk is then the normalised improvement of adversarial success above this baseline:
$\text{Risk}_{j}=\max\left\{0,\frac{TCAP^{\prime}_{j}-WEAP_{j}}{1-WEAP_{j}}\right\}$ (C.6)
A value near 0 indicates synthetic data provides no additional risk beyond baseline, while a value near $1$ signals a maximum increase in disclosure risk. Negative values (adversary performs worse than random) are truncated to zero by convention (ran2024multiobjective; elliot2023samples).
Intuition: This metric isolates the extra “leakage” in matching risk—beyond what already exists in the univariate distribution of the target — arising from access to the synthetic data. It thus accounts for singling out, linkage and inference risks (due to the empirical distribution of sensitive values per equivalence class) identified by  A29WP as being critical determinants of risks to anonymity. The disclosure risk for the dataset is the mean of $R_{j}$ over all synthetic records meeting the $\tau$ threshold. The pseudocode to calculate TCAP can be seen in Algorithm 7
Algorithm 7 TCAP calculation
Input: Original data $\mathcal{D}$, synthetic data $\mathcal{D}^{\prime}$, key variables $K$, target variable $T$, threshold $\tau$
1:for each synthetic record $j$ in $\mathcal{D}^{\prime}$ do
2:  Identify equivalence class $K^{\prime}_{j}$ and $T^{\prime}_{j},K^{\prime}_{j}]$ in synthetic data
3:  $p^{\prime}\leftarrow\frac{\mathbb{I}[T^{\prime}_{j},K^{\prime}_{j}]}{\mathbb{I}[K^{\prime}_{j}]}$
4:  if $p^{\prime}<\tau$ then
5:   Continue to next $j$ $\triangleright$ Skip if class too small
6:  end if
7:  Identify all records in $\mathcal{D}$ with $K_{i}=K^{\prime}_{j}$, denote as class $C$
8:  if $C$ is empty then
9:   $R_{j}\leftarrow 0$, go to next $j$
10:  end if
11:  Calculate $\mathrm{TCAP}^{\prime}_{j}$ using equation C.4
12:  Calculate $\mathrm{WEAP}_{j}$ using equation C.5
13:  Calculate $\text{Risk}_{j}$ using equation C.6
14:end for
[…]
C.3 Reproducibility
This appendix provides the procedural summary sufficient for manual replication or verification by independent researchers. Table 57 showed detailed variables implemented in the utility and risk evaluation.
Table 5: Variables used for ROC evaluation in each dataset.
Dataset ROC Variables (Key/Categorical features)
UK ECONPRIM, ETHGROUP, LTILL, QUALNUM, SEX, SOCLASS, TENURE, MSTATUS
Canada ABIDENT, SEX, TENURE, URBAN, BPLMOM, BPLPOP, CITIZEN, LANG, MARST, RELATE, MINORITY, RELIG, BPL
Fiji PROV, TENURE, RELATE, SEX, ETHNIC, MARST, RELIGION, BPLPROV, RESPROV, RESSTAT, SCHOOL, TRAVEL
Rwanda STATUS, SEX, URBAN, OWNERSH, DISAB2, DISAB1, RELATE, RELIG, HINS, NATION, BPL
Indonesia OWNERSHIP, LANDOWN, RELATE, SEX, MARST, HOMEMALE, RELIGION, SCHOOL, LIT, EDATTAIND, DISABLED
Adult workclass, education, marital-status, occupation, relationship, race, sex, native-country, income
Churn Geography, Gender, Tenure, NumOfProducts, HasCrCard, IsActiveMember, Exited
Table 6: Variables used for CIO evaluation (targets and explanatory variables) in each dataset. Note that for CIO mainly uses logistic regression as the model, considering the categorical target variables, except the CreditScore and EstimatedSalary in Churn that uses linear regression since they are continuous variables.
Dataset Target Variables Explanatory (Key) Variables
UK TENURE, MSTATUS ECONPRIM, ETHGROUP, LTILL, QUALNUM, SEX, SOCLASS, TENURE, MSTATUS, AGE
Canada TENURE, MARST ABIDENT, CLASSWK, DEGREE, EMPSTAT, SEX, URBAN, TENURE, MARST, AGE, HRSWK, INCTOT, WKSWORK
Fiji TENURE, MARST CLASSWKR, ETHNIC, RELIGION, EDATTAIN, SEX, PROV, TENURE, MARST, AGE
Rwanda OWNERSH, MARST DISAB1, EDCERT, CLASSWK, LIT, RELIG, SEX, OWNERSH, MARST, AGE
Indonesia OWNERSHIP, MARST LANDOWN, RELATE, SEX, HOMEFEM, HOMEMALE, RELIGION, LIT, SCHOOL, EDATTAIND, DISABLED, OWNERSHIP, MARST, AGE
Adult income, marital-status workclass, education-num, marital-status, occupation, relationship, race, sex, native-country, income, age, fnlwgt, capital-gain, capital-loss, hours-per-week
Churn Exited, CreditScore, EstimatedSalary Geography, Gender, Tenure, NumOfProducts, HasCrCard, IsActiveMember, Exited, CreditScore, Age, Balance, EstimatedSalary
Table 7: Variables used for TCAP calculation (target and key classes) in each dataset.
Dataset Target Variables Key Variables for Equivalence Class Definition
UK TENURE, MSTATUS ECONPRIM, ETHGROUP, LTILL, QUALNUM, SEX, SOCLASS, TENURE, MSTATUS
Canada TENURE, MARST ABIDENT, CLASSWK, DEGREE, EMPSTAT, SEX, URBAN, TENURE, MARST
Fiji TENURE, MARST CLASSWKR, ETHNIC, RELIGION, EDATTAIN, SEX, PROV, TENURE, MARST
Rwanda OWNERSH, MARST DISAB1, EDCERT, CLASSWK, LIT, RELIG, SEX, OWNERSH, MARST
Indonesia OWNERSHIP, MARST LANDOWN, RELATE, SEX, HOMEFEM, HOMEMALE, RELIGION, LIT, SCHOOL, EDATTAIND, DISABLED, OWNERSHIP, MARST
Adult income, marital-status workclass, education-num, marital-status, occupation, relationship, race, sex, native-country, income
Churn Exited, CreditScore, EstimatedSalary Geography, Gender, Tenure, NumOfProducts, HasCrCard, IsActiveMember, Exited
Appendix D Additional and Detailed Results
D.1 Variance Comparison Results
[…]
Table 8: Comparative performance of synthetic data results using theoretical (left) and relaxed (right) variance for TabbyFlow-OT. The table highlights that relaxing the theoretical variance to learn the distribution of continuous variables frequently leads to better performance.
$0.5A_{t}(x_{t})^{-2}$ $0.5A_{t}(x_{t})^{-1}$ (ours)
dataset Utility Risk Utility Risk
UK 0.5137 $\pm$ 0.0072 0.3998 $\pm$ 0.0074 0.8333 $\pm$ 0.0191 0.4889 $\pm$ 0.0043
Indonesia 0.5149 $\pm$ 0.0085 0.3280 $\pm$ 0.0141 0.6671 $\pm$ 0.0046 0.4114 $\pm$ 0.0149
Canada 0.7479 $\pm$ 0.0109 0.2961 $\pm$ 0.0175 0.7718 $\pm$ 0.0211 0.3206 $\pm$ 0.0146
Fiji 0.7002 $\pm$ 0.0228 0.5574 $\pm$ 0.0060 0.7263 $\pm$ 0.0205 0.5705 $\pm$ 0.0063
Rwanda 0.7319 $\pm$ 0.0199 0.5019 $\pm$ 0.0125 0.7284 $\pm$ 0.0231 0.5379 $\pm$ 0.0139
Adult 0.7374 $\pm$ 0.0155 0.4917 $\pm$ 0.0162 0.7720 $\pm$ 0.0231 0.5443 $\pm$ 0.0177
Churn 0.7613 $\pm$ 0.0103 0.0635 $\pm$ 0.0792 0.7966 $\pm$ 0.0098 0.0773 $\pm$ 0.0579
D.2 Main Comparison on ID, AD, CH Datasets
Table 9 presents a comparative evaluation of six synthetic data generation algorithms across four datasets (UK, ID, AD, CH), highlighting the two best-performing algorithms per dataset. Overall, TabSynFlow-OT consistently achieves superior utility in the ID and CH datasets compared to diffusion baselines, and matches TabSyn while delivering lower disclosure risk in AD. Notably, TabSynFlow-OT also provides reduced risk in ID and CH, making it an effective choice for synthesising tabular data which increases utility and reduces privacy risk.
By contrast, TabbyFlow-OT yields the highest utility for AD, a result that is accompanied by increased disclosure risk, while in ID, it generates synthetic data with lower utility but achieves a substantial reduction in risk. This pattern illustrates how variations in model performance and privacy exposure reflect the empirical relationship between utility and risk across different datasets. TabSyn consistently ranks as the second-best in utility across all three datasets.
By combining with the main text, these results underscore the strength of flow matching approaches for tabular data synthesis, particularly the flexibility of TabSynFlow-OT as an alternative to diffusion-based models in extreme number of observations, i.e. too high or too low.
Table 9: Comparative Performance of six synthetic data generation algorithms across three datasets (ID, AD, CH). Bold and underline indicate the first and second best performing algorithms for each dataset, respectively. The table highlights the comparable performance of TabbyFlow and TabSyn, while TabSynFlow-OT stands out as a strong alternative—particularly in the ID dataset, where TabbyFlow exhibits higher risk in conjunction with generated high utility synthetic data.
Evaluation Algorithm ID AD CH
[…]
D.3 Number of Function Evaluations
Figure 3 shows how computational cost is related to the quality of the output between methods (TabSyn, TabSynFlow and TabbyFlow) and datasets. In general, both utility and risk increase as NFEs grow. In four datasets (UK, RW, CA, FI), TabbyFlow seems to outperform TabSyn. However, in ID and CH, TabSynFlow-OT outperforms TabSyn. Therefore, if one has a limited computational budget, they may choose TabbyFlow or TabSynFlow using the OT path as the algorithm that performed best on low-step setting among all algorithms, particularly TabSyn.
Figure 3: Utility and risk evaluation of TabSynFlow and TabbyFlow based on dataset and average using different number of evaluations with formula $2^{n},n=[2,...,10]$. TabbyFlow-OT (and TabSynFlow-OT) achieve strong utility at low NFEs ($\leq 100$) and tend to converge by $\approx$128, whereas TabSyn requires higher NFEs to catch up—so OT + flow-matching is attractive under tight compute budgets.
D.4 Integration Time
While the original implementation of flow matching uses full ODE integration, it is possible to terminate the ODE at intermediate times—a principle explored in this section. Our analysis of ODE integration times ($t_{ode}$) in Figure 4, spanning datasets including Rwanda, Adult, Churn, UK, Indonesia, Canada, and Fiji, offers critical insights into how utility and risk evolve during synthetic data generation. For TabSynFlow with optimal transport (OT), early stopping at $t_{ode}=0.6$ consistently produces high-utility synthetic data, whereas variance preserving (VP) paths in both TabSynFlow and TabbyFlow tend to underperform at this stage. This indicates that OT trajectories better retain the data structure in earlier integration steps, while VP introduces noise prematurely. Notably, the Rwanda dataset presents an interesting anomaly: TabbyFlow-OT achieves slightly better utility than VP methods at $t_{ode}=0.6$, which may be attributed to the dataset’s relatively simple feature distributions.
Figure 5 showed the utility and risk in details when $0.9\leq t_{ode}\leq 1$. At $t=0.9$, OT methods maintain superior performance, particularly for the UK and Adult datasets, although the gap between OT and VP narrows for the latter. This implies dataset-specific sensitivity to VP’s noise characteristics. Near full integration ($t_{ode}\rightarrow 1.0$), all algorithms exhibit erratic behaviour in some datasets such as Fiji: the utility drops and the risk increases. This problem suggests that flow matching is prone to quality drops when fully integrated.
Figure 4: Effect of ODE integration time on utility–risk (per dataset + average) for TabSynFlow and TabbyFlow. Utility and risk evaluation of TabSynFlow and TabbyFlow based on dataset and average on different integration time from $t_{ode}=[0.6,1]$ with interval 0.1. Colors encode method and path (VP vs OT). OT achieves high-utility solutions early (e.g., $t_{ode}=0.6$), while pushing to $t_{ode}\to 1$ often reduces utility and/or increases risk—so early stopping is preferable.
Figure 5: Utility and risk evaluation of TabSynFlow and TabbyFlow based on dataset and average on late integration time $t=[0.9,0.95,0.975,1]$. For selected datasets, we zoom into the late integration window using the same color/marker scheme as Figure 4. Take-home: Near full integration, instability emerges across several datasets (e.g., Fiji), consistent with utility drops and risk increases; OT remains comparatively robust but still degrades as $t_{ode}\to 1$.
Appendix E Flow Matching Results on Different Interpolants
E.1 Interpolation Path in Flow Matching
A central element in continuous-time generative modelling is the choice of interpolation schedule, which determines how the clean data $x_{1}$ is progressively transformed into noise, and conversely how the noise is mapped back to the data. This schedule is characterised by two functions: the signal decay factor $\alpha(t)$ and the noise scale $\sigma(t)$. Together, these define the marginal distribution.
$x_{t}\sim\mathcal{N}\big(\alpha(t)x_{1},\sigma^{2}(t)I\big),$
as well as their time derivatives $\dot{\alpha}(t)$ and $\dot{\sigma}(t)$, which govern the dynamics in both ODE and SDE formulations.
Different schedules entail distinct trade-offs between training stability, sample quality, and computational efficiency. While the main text provides the primer for the OT and VP paths, flow matching formulations can also use alternative paths such as variance-exploding (lipman2023flowmatchinggenerativemodeling), cosine (albergo2023buildingnormalizingflowsstochastic), or logit-normal (labs2025flux1kontextflowmatching) interpolants. In this section, we summarise these schedules in two comparative tables: Table 10 for the interpolants $(\alpha(t),\sigma(t))$ and Table 11 for their derivatives $(\dot{\alpha}(t),\dot{\sigma}(t))$.
Table 10: Interpolants $\alpha(t)$ and $\sigma(t)$ for different schedules.
Interpolation
$\alpha(t)$ $\sigma(t)$
Linear 1t1-t $t$
Variance-Preserving $\exp\!\Big(-\tfrac{1}{2}\int_{0}^{1-t}\beta(s)\,ds\Big)$ $\sqrt{1-\alpha^{2}(t)}$
Variance-Exploding 11 $\sigma_{\min}\Big(\tfrac{\sigma_{\max}}{\sigma_{\min}}\Big)^{t}$
Cosine $\sin\!\big(\dfrac{\pi}{2}t\big)$ $\cos\!\big(\dfrac{\pi}{2}t\big)$
Logit-Normal
$\dfrac{e^{\mu}}{e^{\mu}+(\tfrac{1}{ct}-1)^{\lambda}}$
$\mu=\log 3,\,\lambda=1$
$1-\alpha(t)$
Table 11: Derivative of interpolants for different schedules.
Interpolation
$\dot{\alpha}(t)$ $\dot{\sigma}(t)$
Linear 1-1 11
Variance-Preserving $\tfrac{1}{2}\beta(t)\alpha(t)$ $\tfrac{1}{2}\dfrac{\beta(t)\alpha^{2}(t)}{\sigma(t)}$
Variance-Exploding 0 $\sigma(t)\,\ln\!\Big(\tfrac{\sigma_{\max}}{\sigma_{\min}}\Big)$
Cosine $\dfrac{\pi}{2}\,\cos\!\big(\dfrac{\pi}{2}t\big)$ $\dfrac{\pi}{2}\,\sin\!\big(\dfrac{\pi}{2}t\big)$
Logit-Normal $\dfrac{\lambda\exp(\mu)}{ct^{2}(e^{\mu}+(\tfrac{1}{ct}-1)^{\lambda})^{2}}\Big(\tfrac{1}{ct}-1\Big)^{\lambda-1}$ $\dot{\alpha}(t)$
E.2 Empirical Results on TabSynFlow
Table 12 disseminated the simulation results of TabSynFlow using different paths. Based on the table, it seems that the comparative analysis shows clear distinctions between the paths. optimal transport (OT) offers the best balance of utility and risk. variance preserving (VP) and Cosine paths reduce risk further, making them well suited for privacy-sensitive contexts, though at a slight cost to utility. In contrast, the variance exploding (VE) and Logit paths underperforms on both measures. Overall, OT is a strong default, VP is preferable when privacy is prioritised, and Cosine remains a viable alternative, although path choice should be validated for each dataset.
Table 12: Comparison of utility and risk across different paths on TabSynFlow.
Dataset OT VP VE Cosine Logit-Normal
Utility Risk Utility Risk Utility Risk Utility Risk Utility Risk
UK 0.7796 0.4834 0.6696 0.4630 0.2387 0.5263 0.7554 0.4802 0.5435 0.4622
Indonesia 0.8981 0.6616 0.7814 0.6464 0.4467 0.6186 0.8732 0.6547 0.5206 0.4782
Canada 0.7035 0.2493 0.6403 0.2073 0.2841 0.3827 0.7021 0.2485 0.4875 0.2233
Fiji 0.6798 0.5506 0.6014 0.5208 0.2800 0.5376 0.7115 0.5526 0.3465 0.4593
Rwanda 0.6765 0.4894 0.5875 0.4655 0.2723 0.5250 0.6681 0.4923 0.4528 0.4513
Adult 0.7560 0.4693 0.6843 0.4131 0.2745 0.5023 0.7466 0.4675 0.4429 0.2489
Churn 0.8784 0.0429 0.8126 0.0246 0.4922 0.0216 0.8284 0.0562 0.7189 0.0630
Average 0.7674 0.4209 0.6824 0.3915 0.3269 0.4449 0.7551 0.4217 0.5018 0.3409
E.3 Empirical Results on TabbyFlow
The results on Table 13 highlights clear differences of the TabbyFlow results between paths. optimal transport (OT) delivers strong overall performance and serves as a reliable default, while Logit performs comparably and in some cases slightly better. VP and Cosine stand out as the least risk, making them suitable when privacy is prioritised, though they sacrifice a little utility. Variance exploding (VE), by contrast, shows worst performance, limiting its general usefulness. These findings highlight that trajectory choice should align with application needs: OT offers balanced performance, VP prioritises privacy protection, while Logit and cosine present viable alternatives to OT and VP, respectively.
Table 13: Comparison of Utility and Risk Across Different Paths on TabbyFlow
Dataset OT VP VE Cosine Logit-Normal
Utility Risk Utility Risk Utility Risk Utility Risk Utility Risk
UK 0.8333 0.4889 0.8066 0.4765 0.3238 0.5331 0.8001 0.4721 0.8201 0.4887
Indonesia 0.6671 0.4114 0.6488 0.3918 0.4519 0.3955 0.6319 0.3856 0.6520 0.3888
Canada 0.7718 0.3206 0.7472 0.3008 0.4083 0.3437 0.7626 0.3102 0.7619 0.3032
Fiji 0.7263 0.5705 0.7451 0.5588 0.3052 0.6000 0.7121 0.5341 0.7303 0.5694
Rwanda 0.7284 0.5379 0.7358 0.5180 0.3267 0.5222 0.7440 0.5138 0.7396 0.5271
Adult 0.7720 0.5443 0.7581 0.5037 0.3097 0.5666 0.7205 0.5107 0.7665 0.5580
Churn 0.7966 0.0773 0.8066 0.0753 0.5563 0.0481 0.7913 0.0841 0.7982 0.0767
Average 0.7565 0.4215 0.7498 0.4036 0.3831 0.4299 0.7375 0.4015 0.7526 0.4160
E.4 SDE Performance on Different $g_{t}$
Tables 14 and 15 disseminate the comparison of the difference in utility ($\Delta\text{U}$) and risk ($\Delta\text{R}$) when using SDE with different $g_{t}$ on TabbyFlow compared to its ODE counterparts as a baseline (See Table 13). The tables reveal consistently small deviations across all datasets, with the majority of differences within [0.02,+0.02][-0.02,+0.02]. This minimal variation strongly validates the theoretical equivalence between the SDE and ODE marginals, demonstrating that SDEs preserve the same distributional properties regardless of the specific $g_{t}$ function employed. The results confirm that the controlled stochasticity introduced through SDE sampling maintains distributional fidelity while offering additional flexibility in the sampling process.
In particular, several configurations show the desirable outcome of simultaneously improving both utility and reducing risk. For example, Rwanda in the OT configuration achieves $\Delta\text{U}=+0.0105$ and $\Delta\text{R}=-0.0054$, while Fiji in the VP configuration achieves $\Delta\text{U}=+0.0067$ and $\Delta\text{R}=-0.0017$. These dual improvements indicate that SDE-based approaches can not only match the performance of ODE but, in some cases, produced high utility data with lower privacy risk. This dataset-dependent behavior underscores the importance of empirical selection for $g_{t}$ in synthesiser design, as the configuration can vary based on feature complexity, data distribution, or presence of rare categories.
Table 14: Difference of utility and risk across different $g_{t}$ setting in SDE w.r.t. the ODE results on TabbyFlow-OT.
Dataset OT VP VE Cos Logit
$\Delta\text{U}(\uparrow)$ $\Delta\text{R}(\downarrow)$ $\Delta\text{U}(\uparrow)$ $\Delta\text{R}(\downarrow)$ $\Delta\text{U}(\uparrow)$ $\Delta\text{R}(\downarrow)$ $\Delta\text{U}(\uparrow)$ $\Delta\text{R}(\downarrow)$ $\Delta\text{U}(\uparrow)$ $\Delta\text{R}(\downarrow)$
UK -0.0073 -0.0011 -0.0044 0.0000 -0.0066 -0.0007 -0.0024 0.0003 -0.0103 0.0022
Indonesia -0.0044 -0.0130 -0.0105 -0.0106 -0.0028 -0.0072 -0.0075 -0.0041 -0.0027 -0.0005
Canada -0.0011 -0.0020 0.0060 0.0007 0.0004 -0.0026 0.0047 -0.0010 0.0034 -0.0076
Fiji 0.0006 -0.0024 0.0067 -0.0017 0.0042 -0.0027 0.0010 -0.0052 0.0029 -0.0001
Rwanda 0.0105 -0.0054 0.0142 0.0004 0.0093 -0.0050 0.0123 -0.0039 0.0112 -0.0023
Adult 0.0077 0.0012 0.0027 0.0040 0.0047 0.0021 0.0067 0.0038 0.0039 0.0045
Churn 0.0036 -0.0156 0.0021 0.0070 0.0025 -0.0047 0.0024 0.0112 0.0032 0.0210
Table 15: Difference of utility and risk across different $g_{t}$ setting in SDE w.r.t. the ODE results on TabbyFlow-VP.
Dataset OT VP VE Cos Logit
$\Delta\text{U}(\uparrow)$ $\Delta\text{R}(\downarrow)$ $\Delta\text{U}(\uparrow)$ $\Delta\text{R}(\downarrow)$ $\Delta\text{U}(\uparrow)$ $\Delta\text{R}(\downarrow)$ $\Delta\text{U}(\uparrow)$ $\Delta\text{R}(\downarrow)$ $\Delta\text{U}(\uparrow)$ $\Delta\text{R}(\downarrow)$
UK -0.0035 -0.0005 -0.0070 -0.0024 -0.0101 -0.0020 -0.0072 -0.0022 -0.0045 -0.0010
Indonesia -0.0034 0.0028 -0.0069 0.0001 -0.0025 0.0077 -0.0030 0.0023 -0.0034 0.0017
Canada 0.0015 0.0029 0.0025 -0.0027 0.0014 0.0017 0.0026 0.0007 0.0048 0.0076
Fiji 0.0038 -0.0012 0.0011 -0.0017 0.0097 -0.0022 0.0021 -0.0026 0.0095 -0.0006
Rwanda 0.0103 0.0012 0.0038 0.0000 0.0032 -0.0030 0.0084 0.0054 0.0098 -0.0017
Adult 0.0043 -0.0035 0.0082 -0.0021 -0.0006 -0.0028 0.0055 0.0010 0.0004 0.0006
Churn 0.0001 -0.0102 0.0002 0.0018 0.0009 -0.0014 -0.0009 0.0061 -0.0011 -0.0022